# M. Hashem Pesaran

### v1

@@ Abstract
Forecasts play a central role in decision making under uncertainty. After a brief review of the general issues, this paper considers ways of using high-dimensional data in forecasting. We consider selecting variables from a known active set, known knowns, using Lasso and OCMT, and approximating unobserved latent factors, known unknowns, by various means. This combines both sparse and dense approaches. We demonstrate the various issues involved in variable selection in a high-dimensional setting with an application to forecasting UK inflation at different horizons over the period 2020q1-2023q1. This application shows both the power of parsimonious models and the importance of allowing for global variables.
@@ 1 Introduction
If both statistical and policy models are used, there is an issue as to how to integrate them. In September 2018, the Bank of England Independent Evaluation Office, IEO reported back to Court of the Bank on the implementation and impact of the 2015 IEO review of forecasting performance. In response some ’non-structural’ models had been introduced as a source of challenge, and outputs were routinely shown to the Monetary Policy Committee as a way of cross checking the main forecast. Some but not all members found them helpful, but there was no desire to develop more models of this sort. Internally they had not been integrated into the forecast process as a source of challenge.” Forecast averaging has been widely shown to improve forecast performance, but forming averages on many variables may lack coherence and consistency.
@@ 2 The high-dimensional forecasting problem
To be more specific, ignoring intercepts and factors which we introduce below, suppose $y_{t}$ is the first element of the high-diemnsional vector $\boldsymbol{w}_{t}$, assumed to follow the first order VAR model,
@@
where $\boldsymbol{\phi}_{h}^{\prime}$ is the first row of $\boldsymbol{\Phi}^{h},$ and $u_{h,t+h}$ is the first element of $\mathbf{u}_{h,t+h}$. We still face a high-dimensional problem since there are a large number of potential covariates in $\boldsymbol{w}_{t}$. We consider the implementation of the direct approach under two scenarios concerning the potential predictors. First, when it is known that the target variable $y_{t+h}$ is a sparse linear function of a large set of observed variables $\boldsymbol{x}_{t}$ (a subset of $\boldsymbol{w}_{t}$) in the sense that $y_{t+h}$ depends on a small number of covariates that are known to be a subset of $\boldsymbol{w}_{t},$ known as the ‘active set’. The machine learning literature focuses on this case, which we refer to as the case of ”known knowns”. Second, when $y_{t+h}$ could also depend on a few latent (unobserved) factors, $\mathbf{f}_{t}$, not directly included in the active set, which we call the case of ”known unknowns”.
@@
for $t=1,2,...,T-h$, where $c_{h}$ is the intercept, $\mathbf{z}_{t}$ is a vector of small number, $p$, of pre-selected covariates included across all horizons $h$. Obvious examples, include lagged values of the target variable $(y_{t},y_{t-1},...)$. Other variables can also be included in $\mathbf{z}_{t}$ on the basis of a priori theory or strong beliefs. The second component of $y_{t+h}$ specifies the subset of variables in the active set $\boldsymbol{x}_{t}=\left(x_{1t},x_{2t},....,x_{Kt}\right)^{\prime}$. $I(j\in DGP)$ is an indicator variable which takes the value of unity if $x_{jt}$ is included in the data generating process (DGP) for $y_{t+h}$ and zero otherwise. It is only if $I(j\in DGP)=1$ that $\beta_{jh}$ will be identified. The number of variables included in the DGP is given by $k=\sum_{j=1}^{K}I(j\in DGP)$, which is supposed to be small and fixed as $T$ (and possibly $K$) become large. This assumption imposes sparsity on the relationship between the target and the variables in the active set. In addition, we allow for a small number of latent factors, $\mathbf{f}_{t},$ that represent other variables influencing $y_{t+h}$ that are not observed directly, but known to be present - the known unknowns.
Giannone, Lenza, and Primiceri (2021) contrast sparse methods, that select a few variables from the active set as predictors, such as Lasso and OCMT discussed below, and dense methods, that select all the variables in the active set but attach small weights to many of them, such as principal components (PC), ridge regression and other shrinkage techniques. Rather than having to choose between sparse and dense predictors, we consider approaches that combine the two. We apply sparse selection methods to the variables in the active set, and use dense shrinkage methods to approximate $\mathbf{f}_{t}$ from a wider set of variables with $\boldsymbol{x}_{t}$ included as a subset. We first consider the selection problem, known knowns, where we know the active set of potential covariates, and we then consider known unknowns where there are unobserved factors.
@@ 3 Known knowns
subject to some penalty condition on $\{\beta_{hj}\}$. Some of the covariates, $x_{jt}$, could be transformations of other covariates, such as interaction terms. It is assumed that the model is correctly specified, in the sense that, apart from $\boldsymbol{z}_{t}$, the variables that drive $y_{t+h}$ are all included in the active set, $\boldsymbol{x}_{t}=\left(x_{1t},x_{2t},....,x_{Kt}\right)^{\prime}$.
Many standard forecasting techniques result from the particular choice of the penalty function. Shrinkage estimators such as ridge or some Bayesian forecasts can be derived using the $\ell_{2}$ norm $\sum_{i=1}^{K}$ $\beta_{hj}^{2}<C_{h}<\infty.$ Lasso (least absolute shrinkage and selection operator) follows when the $\ell_{1}$ norm is used $\sum_{i=1}^{K}\left|\beta_{hj}\right|<C_{h}<\infty.$ The difference is shown in Figure 1 below in Tibshirani (1996) where the $\ell_{1}$ norm yields corner solutions with many of the coefficients, $\beta_{hj}$, estimated to be zero. In contrast, the use of $\ell_{2}$ norm yields non-zero estimates for all the coefficients with many very close to zero.
@@
There are a very large number of penalized techniques. Elastic net regularisation has both an $\ell_{1}$ and an $\ell_{2}$ penalty, with ridge regression and Lasso being special cases. There are also a large number of variants of Lasso, including adaptive Lasso, grouped Lasso, double Lasso. We will focus on Lasso itself, which we use in our empirical application and which we will compare to OCMT as an alternative procedure which is based on inferential rather than penalized procedures.
@@ 3.1 Lasso
Lasso estimates $\boldsymbol{\beta}_{h}$ by solving the following optimization problem:
@@ 3.2 OCMT
The need for cross-validation is avoided in the procedure proposed by Chudik, Kapetanios, and Pesaran (2018), (CKP). This is the ”One Covariate at a time Multiple Testing” (OCMT) procedure, where covariates are selected one at a time, using $t-$statistic for testing the significance of the variables in the active set, individually. [^5] Since $t$-ratios are invariant to scale, no pre-standardization of the covariates in the active set is required. This is in contrast to Lasso which is typically implemented after in-sample standardization of the covariates. Ideas from the multiple testing literature are used to control the false discovery rate, and ensure the selected covariates encompass the true covariates (signals) with probability tending to unity, under certain regularity conditions. Like Lasso, OCMT has no difficulty in dealing with (pure) noise variables, and is very effective at eliminating them. Also, like Lasso, it requires some $min$ condition such as $\left|\beta_{jh}^{0}\right|>>\sqrt{\frac{k\log(K)}{T}}$, for $j=1,2,...,k$ [^6] To simplify the exposition we have dropped explicit reference to the forecast horizon, $h$. But in practice, and as we shall see from the empirical applications below, the number and identity of selected signals could differ with $h$. But because it considers a single variable at a time, OCMT does not require the IRC condition to hold and is not affected by the correlation between the members of the DGP as Lasso is. Instead it requires the number of pseudo signals, say $k_{T}^{\ast}$, to rise not faster than $\sqrt{T}$.  Chudik, Pesaran, and Sharifvaghefi (2023), discussed below, is primarily concerned with parameter instability, but Section 4 of that paper has a detailed comparison of the assumptions required for Lasso and OCMT under parameter stability.
@@
where as before $\boldsymbol{z}_{t}$ is a known vector of pre-selected variables, and it is assumed that the $k$ signals are contained in the known active set $\mathcal{S}_{K,t}=\left\{x_{jt},j=1,2,...,K\right\}$. Note that for now the DGP in (7) does not include the additional latent factors, $\mathbf{f}_{t}$, introduced in (3). Without loss of generality, consider the extreme case where there are no noise variables and all proxy or pseudo signal variables ($x_{jt}$, for $j=k+1,k+2,...,K)$ are correlated with the signals, $\mathbf{x}_{1t}=(x_{1t,}x_{2t},...,x_{kt})^{\prime}$. In this case $k_{T}^{\ast}$ rises with $K$ and OCMT is no longer applicable. However, in this case by construction the signals, $\mathbf{x}_{1t}$, become latent factors for the proxy variables and we have
@@
for $j=k+1,k+2,....,K$. Although the identity of these common factors are unknown, they can be approximated by the principal components of the variables in the active set. Specifically, following Sharifvaghefi (2023), denote the latent factors that result in non-zero correlations between the noise variables in the active set and the signals by $\boldsymbol{\varkappa}_{t}$ and consider the factor model
@@
where $\mathbf{b}_{h}=\sum_{j=1}^{K}I(j\in DGP)\beta_{jh}\boldsymbol{\kappa}_{j}$. When $\boldsymbol{\varkappa}_{t}$ and $v_{jt}$ are known, the problem reduces to selecting $v_{jt}$ from $\mathcal{S}_{K,t}^{v}=\left\{v_{jt},j=1,2,...,K\right\},$ conditional on $\mathbf{z}_{t}$ and $\boldsymbol{\varkappa}_{t}$. Sharifvaghefi shows that the OCMT selection can be carried out using the principal component estimators of $\boldsymbol{\varkappa}_{t}$ and $\mathbf{v}_{jt}$ denoted by $\boldsymbol{\hat{\varkappa}}_{t}$ and $\mathbf{\hat{v}}_{jt}$, if both $K$ and $T$ are large. He labels this procedure as generalized OCMT (GOCMT).
@@ 3.3 GOCMT
$p$ is the nominal size (usually set to $5\%$), $\Phi^{-1}(\cdot)$ is the inverse of a standard normal distribution function and $\delta$ is a fixed constant set in the interval $[1,1.5]$. In the second step a multivariate regression of $y_{t+h}$ on $\boldsymbol{z}_{t}$ and all the selected regressors is considered for inference and forecasting. Note that at the moment it is assumed that $\boldsymbol{\varkappa}_{t}$ does not directly affect $y_{t+h},$ it only enters through the $x_{jt}.$ It represents the signals, the common factors correlated with the proxies, and provides a way of filtering the correlations in the first step.
When the covariates are not highly correlated, OCMT applies irrespective of whether $K$ is small or large relative to $T$, so long as $T=\ominus\left(K^{c}\right)$, for any finite $c>0$.  But to allow for highly correlated covariates, GOCMT requires $K$ to be sufficiently large to enable the identification of $\boldsymbol{\varkappa}_{t}$. In cases where $K$ is not that large, it might be a good idea to augment the active set for the target variable, $y_{t+h}$, $\mathcal{S}_{K,t}=\left\{x_{jt},\text{ }j=1,2,...,K\right\}$, with covariates for other variables determined simultaneously with $y_{t+h}$, ending up with $\overline{\mathcal{K}}>K$ covariates for identification of $\boldsymbol{\varkappa}_{t}$. GOCMT does not impose any restriction on the correlations between the variables other than that they cannot be perfectly collinear.
@@ 4 Known unknowns
The forecasting problem now involves both selection and shrinkage. The $\mathbf{f}_{t}$ can be identified by, for instance, the PCs of the augmented active set $x_{jt},$ $j=1,2...,K,K+1,...,\overline{\mathcal{K}}$ which can be wider than the active set of covariates used to predict $y_{t}$. There is a variety of other ways that the unobserved $\mathbf{f}_{t}$ could be estimated, but we use PCs as an example, since they are widely used.
The missing factors are unlikely to be only specific to the target variable under consideration. For example, if $y_{t+h}$ is UK inflation, the missing factor(s) could be external to the UK. Inflation and output growth in other countries could be equally relevant. Call this vector of global factors $\mathbf{g}_{t}$. A natural extension is to introduce forecast equations for other countries (entities) who have close trading relationships with UK and use penalized panel regressions, where the panel dimension allows identification of the known unknowns.
Suppose there are $N$ other units (countries) that are affected by observed country specific covariates, $\mathbf{z}_{it},$ $i=1,2,...,N,$ and $x_{i,jt}$ for $j=1,2,...,K_{i},$ plus domestic latent factors $\mathbf{f}_{it}$, and global latent factors, $\mathbf{g}_{t}$. The forecasting equations are now generalized as
@@
where $k_{i}=\sum_{j=1}^{K_{i}}I(j\in DGP_{i})$ is small with $K_{i}$ quite large, for $i=0,1,2,...,n$. For the country-specific covariates we postulate that there is an augmented active set

### v2

@@ Abstract
Forecasts play a central role in decision making under uncertainty. After a brief review of the general issues, this paper considers ways of using high-dimensional data in forecasting. We consider selecting variables from a known active set, known knowns, using Lasso and OCMT, and approximating unobserved latent factors, known unknowns, by various means. This combines both sparse and dense approaches to forecasting. We demonstrate the various issues involved in variable selection in a high-dimensional setting with an application to forecasting UK inflation at different horizons over the period 2020q1-2023q1. This application shows both the power of parsimonious models and the importance of allowing for global variables.
@@ 1 Introduction
If both statistical and policy models are used, there is an issue as to how to integrate them. Forecast averaging has been widely shown to improve forecast performance, but forming averages on many variables may lack coherence and consistency. In September 2018, the Bank of England Independent Evaluation Office, IEO reported back to Court of the Bank on the implementation and impact of the 2015 IEO review of forecasting performance. IEO reported that some ’non-structural’ models had been introduced as a source of challenge, and outputs were routinely shown to the Monetary Policy Committee as a way of cross checking the main forecast. While, some but not all members found them helpful, but there was no desire to develop more models of this sort and internally they had not been integrated into the forecast process as a source of challenge.
@@ 2 The high-dimensional forecasting problem
To be more specific, ignoring intercepts and factors which we introduce below, suppose $y_{t}$ is the first element of the high-dimensional vector $\boldsymbol{w}_{t}$, assumed to follow the first order VAR model,
@@
where $\boldsymbol{\phi}_{h}^{\prime}$ is the first row of $\boldsymbol{\Phi}^{h},$ and $u_{h,t+h}$ is the first element of $\mathbf{u}_{h,t+h}$. We still face a high-dimensional problem since there are a large number of potential covariates in $\boldsymbol{w}_{t}$. We consider the implementation of the direct approach under two scenarios concerning the potential predictors. First, when it is known that the target variable $y_{t+h}$ is a sparse linear function of a large set of observed variables $\boldsymbol{x}_{t}$ (a subset of $\boldsymbol{w}_{t}$) known as the ‘active set’. The model is sparse in the sense that $y_{t+h}$ depends on a small number of covariates, that are known to be a subset of the much larger active set. The machine learning literature focuses on this case, which we refer to as the case of ”known knowns”. Second, when $y_{t+h}$ could also depend on a few latent (unobserved) factors, $\mathbf{f}_{t}$, not directly included in the active set, which we call the case of ”known unknowns”.
@@
for $t=1,2,...,T-h$, where $c_{h}$ is the intercept, $\mathbf{z}_{t}$ is a vector of small number, $p$, of pre-selected covariates included across all horizons $h$. Obvious examples, include lagged values of the target variable $(y_{t},y_{t-1},...)$. Other variables can also be included in $\mathbf{z}_{t}$ on the basis of a priori theory or strong beliefs. The third component of $y_{t+h}$ specifies the subset of variables in the active set $\boldsymbol{x}_{Kt}=\left(x_{1t},x_{2t},....,x_{Kt}\right)^{\prime}$. $I(j\in DGP)$ is an indicator variable which takes the value of unity if $x_{jt}$ is included in the data generating process (DGP) for $y_{t+h}$ and zero otherwise. It is only if $I(j\in DGP)=1$ that $\beta_{jh}$ will be identified. We discuss ways to determine the selection indicator $I(j\in DGP)$ below. The number of variables included in the DGP is given by $k=\sum_{j=1}^{K}I(j\in DGP)$, which is supposed to be small and fixed as $T$ (and possibly $K$) become large. This assumption imposes sparsity on the relationship between the target and the variables in the active set. In addition, we allow for a small number of latent factors, $\mathbf{f}_{t},$ that represent other variables influencing $y_{t+h}$ that are not observed directly, but known to be present - the known unknowns.
Giannone, Lenza, and Primiceri (2021) contrast sparse methods, that select a few variables from the active set as predictors, such as Lasso and OCMT discussed below, and dense methods, that select all the variables in the active set but attach small weights to many of them, such as principal components (PC), ridge regression and other shrinkage techniques. Rather than having to choose between sparse and dense predictors, we consider approaches that combine the two. We apply sparse selection methods to the variables in the active set, and use dense shrinkage methods to approximate $\mathbf{f}_{t}$ from a wider set of variables with $\boldsymbol{x}_{t}$ included as a subset. We first consider the selection problem, known knowns, where we know the active set of potential covariates, and we then consider known unknowns where there are unobserved factors. The elastic net regression of Zou and Hastie (2005) discussed below also combines sparse and dense techniques.
Throughout, we shall assume that the errors, $u_{h,t+h\text{ }}$, in (3) satisfy the orthogonality condition $E\left(u_{h,t+h\text{ }}\left|\mathbf{z}_{t},\boldsymbol{x}_{t},\mathbf{f}_{t}\right.\right)=0,$ for $h\geq 1$. In the context of the high-dimensional VAR model discussed above, this orthogonality condition holds so long as the underlying errors, $\mathbf{u}_{t}$, are serially uncorrelated. This is so despite the fact that due to the use of overlapping observations $u_{h,t+h\text{ }}$ will be serially correlated when $h>1$. This is an important consideration when high-dimensional techniques are applied to select predictors for multi-step ahead forecasting; an issue to which we will return.
@@ 3 Known knowns
subject to some penalty condition on $\{\beta_{hj}\}$. Some of the covariates, $x_{jt}$, could be transformations of other covariates, such as interaction terms. It is assumed that the model is correctly specified, in the sense that, apart from $\boldsymbol{z}_{t}$, the variables that drive $y_{t+h}$ are all included in the active set, $\mathbf{x}_{Kt}$.
Penalized regressions estimate $\boldsymbol{\beta}$ by solving the following optimization problem:
$\min_{\mathbf{a}_{h},\mathbf{\beta}_{h}}\left\{{\displaystyle\sum\limits_{t=1}^{T}}\left(y_{t+h}-\mathbf{a}_{h}^{\prime}\boldsymbol{z}_{t}-\boldsymbol{\beta}_{h}^{\prime}\mathbf{x}_{Kt}\right)^{2}+\lambda_{hT}{\displaystyle\sum\limits_{i=1}^{K}}\left[(1-\alpha){\left|\beta_{hi}\right|}+\alpha\beta_{hi}^{2}\right]\right\},$
where $\boldsymbol{\beta}_{h}=(\beta_{h1},\beta_{h2},...,\beta_{hK})^{\prime}$, for given values of the ”tuning” parameters $\lambda$ and $\alpha$. When $\alpha=1,$ we have ridge regression. When $\alpha=0$, $\lambda_{hT}\neq 0,$ we have the Lasso regression, which is better suited for variable selection. When $\lambda_{hT}\neq 0$ and $\alpha\neq 0$ we have the Zou and Hastie (2005) elastic net regression, which also mixes sparse and dense approaches.
Many standard forecasting techniques result from the particular choice of the penalty function. Shrinkage estimators such as ridge or some Bayesian forecasts can be derived using the $\ell_{2}$ norm $\sum_{i=1}^{K}$ $\beta_{hi}^{2}<C_{h}<\infty.$ Lasso (least absolute shrinkage and selection operator) follows when the $\ell_{1}$ norm is used $\sum_{i=1}^{K}\left|\beta_{hi}\right|<C_{h}<\infty.$ The difference is shown in Figure 1 below in Tibshirani (1996) where the $\ell_{1}$ norm yields corner solutions with many of the coefficients, $\beta_{hj}$, estimated to be zero. In contrast, the use of $\ell_{2}$ norm yields non-zero estimates for all the coefficients with many very close to zero.
@@
There are also a large number of variants of Lasso, including adaptive Lasso, group Lasso, double Lasso, fused Lasso and prior Lasso. We will focus on Lasso itself, which we use in our empirical application and which we will compare to OCMT as an alternative procedure which is based on inferential rather than penalized procedures.
@@ 3.1 Lasso
In this paper we focus on Lasso, but acknowledge that are many variations on Lasso such as: adaptive Lasso, group Lasso, fused Lasso, and prior Lasso. Lasso estimates $\boldsymbol{\beta}_{h}$ by solving the following optimization problem:
@@ 3.2 OCMT
The need for cross-validation is avoided in the procedure proposed by Chudik, Kapetanios, and Pesaran (2018), (CKP). This is the ”One Covariate at a time Multiple Testing” (OCMT) procedure, where covariates are selected one at a time, using $t-$statistic for testing the significance of the variables in the active set, individually. [^5] Since $t$-ratios are invariant to scale, no pre-standardization of the covariates in the active set is required. This is in contrast to Lasso which is typically implemented after in-sample standardization of the covariates. Ideas from the multiple testing literature are used to control the false discovery rate, and ensure the selected covariates encompass the true covariates (signals) with probability tending to unity, under certain regularity conditions. Like Lasso, OCMT has no difficulty in dealing with (pure) noise variables, and is very effective at eliminating them. Also, like Lasso, it requires some $min$ condition such as $\left|\beta_{jh}^{0}\right|>>\sqrt{\frac{k\log(K)}{T}}$, for $j=1,2,...,k$ [^6] To simplify the exposition we have dropped explicit reference to the forecast horizon, $h$. But in practice, and as we shall see from the empirical applications below, the number and identity of selected signals could differ with $h$. But because it considers a single variable at a time, OCMT does not require the IRC condition to hold and is not affected by the correlation between the members of the DGP as Lasso is. Instead it requires the number of proxies signals, say $k_{T}^{\ast}$, to rise no faster than $\sqrt{T}$.  Chudik, Pesaran, and Sharifvaghefi (2023), discussed below, is primarily concerned with parameter instability, but Section 4 of that paper has a detailed comparison of the assumptions required for Lasso and OCMT under parameter stability.
@@
where as before $\boldsymbol{z}_{t}$ is a known vector of pre-selected variables, and it is assumed that the $k$ signals are contained in the known active set $\mathcal{S}_{K,t}=\left\{x_{jt},j=1,2,...,K\right\}$. Note that for now the DGP in (7) does not include the additional latent factors, $\mathbf{f}_{t}$, introduced in (3). Without loss of generality, consider the extreme case where there are no noise variables and all proxy or pseudo signal variables ($x_{jt}$, for $j=k+1,k+2,...,K)$ are correlated with the signals, $\mathbf{x}_{1t}=(x_{1t,}x_{2t},...,x_{kt})^{\prime}$. In this case $k_{T}^{\ast}$ rises with $K$ and OCMT is no longer applicable. However, in this case because of the correlation with the proxies, the signals, $\mathbf{x}_{1t}$, become latent factors for the proxy variables and we have
@@
for $j=k+1,k+2,....,K$. Although the identity of these common factors are unknown, because we do not know the true signals, they can be approximated by the principal components of the variables in the active set.
Specifically, following Sharifvaghefi (2023), denote the latent factors that result in non-zero correlations between the noise variables in the active set and the signals by $\boldsymbol{\varkappa}_{t}$ and consider the factor model
@@
where $\mathbf{b}_{h}=\sum_{j=1}^{K}I(j\in DGP)\beta_{jh}\boldsymbol{\kappa}_{j}$. When $\boldsymbol{\varkappa}_{t}$ and $v_{jt}$ are known, the problem reduces to selecting $v_{jt}$ from $\mathcal{S}_{K,t}^{v}=\left\{v_{jt},j=1,2,...,K\right\},$ conditional on $\mathbf{z}_{t}$ and $\boldsymbol{\varkappa}_{t}$. Sharifvaghefi shows that the OCMT selection can be carried out using the principal component estimators of $\boldsymbol{\varkappa}_{t}$ and $\mathbf{v}_{jt}$ denoted by $\boldsymbol{\hat{\varkappa}}_{t}$ and $\mathbf{\hat{v}}_{jt}$, if both $K$ and $T$ are large. He labels this procedure as generalized OCMT (GOCMT). Note that at the moment it is assumed that $\boldsymbol{\varkappa}_{t}$ does not directly affect $y_{t+h},$ it only enters through the $x_{jt}.$ It represents the signals, the common factors correlated with the proxies, and provides a way of filtering the correlations in the first step.
@@ 3.3 GOCMT
$p$ is the nominal size (usually set to $5\%$), $\Phi^{-1}(\cdot)$ is the inverse of a standard normal distribution function and $\delta$ is a fixed constant set in the interval $[1,1.5]$. In the second step a multivariate regression of $y_{t+h}$ on $\boldsymbol{z}_{t}$ and all the selected regressors is considered for inference and forecasting. Serial correlation will arise with OCMT when selection is based on one variable at the time, and the omitted variables are mixing (serially correlated). CKP discuss this in section C of the online theory supplement to their paper and suggest using a more conservative (higher) critical value – namely using $\delta=1.5$ rather than $\delta=1.5$.
When the covariates are not highly correlated, OCMT applies irrespective of whether $K$ is small or large relative to $T$, so long as $T=\ominus\left(K^{c}\right)$, for some finite $c>0$.  But to allow for highly correlated covariates, GOCMT requires $K$ to be sufficiently large to enable the identification of the latent factor, $\boldsymbol{\varkappa}_{t}$. In cases where $K$ is not that large, it might be a good idea to augment the active set for the target variable, $y_{t+h}$, $\mathcal{S}_{K,t}=\left\{x_{jt},\text{ }j=1,2,...,K\right\}$, with covariates for other variables determined simultaneously with $y_{t+h}$, ending up with $\overline{\mathcal{K}}>K$ covariates for identification of $\boldsymbol{\varkappa}_{t}$. GOCMT does not impose any restriction on the correlations between the variables other than that they cannot be perfectly collinear.
@@ 4 Known unknowns
The forecasting problem now involves both selection and shrinkage. The $\mathbf{f}_{t}$ can be identified by, for instance, the PCs of the augmented active set $x_{jt},$ $j=1,2...,K,K+1,...,\overline{\mathcal{K}}$ which can be wider than the active set of covariates used to predict $y_{t}$. There are various other ways that the unobserved $\mathbf{f}_{t}$ could be estimated, but we use PCs as an example, since they are widely used.
The latent factors are unlikely to be only specific to the target variable under consideration. Observed global factors, such as oil and raw material prices or inflation and output growth of major countries such as US can be included in the active set. The main issue is how to deal with global factors, such as technology, political change and so on that are unobserved and tend to affect many countries in the world economy. Call this vector of global factors $\mathbf{g}_{t}$. A natural extension is to introduce forecast equations for other countries (entities) who have close trading relationships with UK and use penalized panel regressions, where the panel dimension allows identification of the known unknowns.
More specifically, suppose there are $N$ other units (countries) that are affected by observed country specific covariates, $\mathbf{z}_{it},$ $i=1,2,...,N,$ and $x_{i,jt}$ for $j=1,2,...,K_{i},$ plus domestic latent factors $\mathbf{f}_{it}$, and global latent factors, $\mathbf{g}_{t}$. The forecasting equations are now generalized as
@@
where $k_{i}=\sum_{j=1}^{K_{i}}I(j\in DGP_{i})$ is finite as $K_{i}\rightarrow\infty$, for $i=0,1,2,...,N$. For the country-specific covariates we postulate that there is an augmented active set